A lunar landing module is descending to the moon’s surface at a steady velocity of 10 m/s. At a height of 120 m, a small object falls from its landing gear. Taking the moon’s gravitational acceleration as 1.6 m/s^2 , at what speed, in m/s does the object strike the moon?
1
Expert's answer
2020-11-17T07:12:49-0500

"\text{let us throw the material point up}"

"h_{max}=\frac{V_0^2}{2a}"

"V_0 =\sqrt{h_{max}*2*a}=\sqrt{120*2*1.6}\approx19.6"

"\text{Upon reaching the surface, the body has the same speed}"

"\text{Add lander speed and get the answer } V=10+19.6 =29.6"

Answer:  "V=29.6\frac{m}{s}"